OpenStudy (anonymous):

how would i solve this using the fundamental theorem of algebra? x^3-10x^2+21x -108 = 0

OpenStudy (phi):

The theorem tells you that there are 3 factors See http://www.mathsisfun.com/algebra/fundamental-theorem-algebra.html you would factor this polynomial. cubics are difficult.... but in this case, try (x-9) divide (x-9) into the polynomial to factor it.

OpenStudy (anonymous):



OpenStudy (phi):

yes,that will work

OpenStudy (anonymous):

what would be the Rational Root Theorem?

OpenStudy (phi):

See http://www.mathwords.com/r/rational_root_theorem.htm but as you see, 108 will give lots of possibilities...

OpenStudy (anonymous):

how would i know which one is q and p?

OpenStudy (phi):

p is one of the factors of the last number -108 and q is the number in front of x^3 (1, which is not usually written because 1*x^3 is x^3) your possible roots will be ± p/q. q is 1, so just ±p now you list all the numbers that divide into 108: 1,2,3, 4, 6, 9, 12, .... (there are lots of them)

OpenStudy (phi):

next time, pick a smaller box!

OpenStudy (anonymous):

well, i thought is was pretty small, i didnt even use a box cuz i wanted smaller, so i used a book

OpenStudy (anonymous):

i have to list ALL #´s that divide into 108?

OpenStudy (phi):

that is what the theorem says... that the root you want is among those numbers. Of course 9 is listed, so it works, but all the other numbers are not roots. If you tested each one, smallest to biggest (by dividing x-a where a is one of the numbers and x+a (could be either + or -) ) you could say, after finding 9, I don't need the rest of the factors of 108....

OpenStudy (phi):

Or you could say you are like the rainman and can do these things in your head.. it's 9!

OpenStudy (anonymous):

i got it, thanks!

OpenStudy (anonymous):



OpenStudy (phi):

one way to find the roots is start with your list ±1,±2,±3, ±4, ±6, ±9, ±12, .... replace x with each possible root: for example x= -1 x^3-10x^2+21x -108 = 0 -1 -10 -21 -108 ≠ 0 and see if you get zero. no with this one with x=1: 1-10+21-180 No keep going until you get to x=+9 729-810+189-108=0 pay dirt once you know 9 is a root, you can divide (as you did up above) to get x^2-x+12=0 use the quadratic formula to show that you do not get any real answers. so x=9 is the only real solution.